OpenStudy (anonymous):

sin θ cos 3θ + cos θ sin 3θ = 0 Use an Addition or Subtraction Formula to simplify the equation.

OpenStudy (anonymous):

Using the Sine addition formula I get Sin4θ, but when I enter that as an answer my homework site marks it as incorrect.

OpenStudy (anonymous):

No, it's as it's shown.

OpenStudy (anonymous):

There are no powers in this problem.

OpenStudy (anonymous):

Consider the left hand side. sin θ cos 3θ + cos θ sin 3θ = sin (θ+3θ) = sin 4θ Do you need to solve it? If not, then I'm pretty sure it is correct. Perhaps you should type s in Sin without capitalising it

OpenStudy (anonymous):

It just asks for me to simplify the equation using the Addition or substraction formula.

OpenStudy (anonymous):

Is there a way to derive Sin4θ into something else?

OpenStudy (anonymous):

Yes. Use double angle formula sin4θ = 2sin2θcos2θ = 2(2sinθcosθ)cos2θ = 4sinθcosθcos2θ I'm pretty sure that would be something undesired

OpenStudy (australopithecus):

sin(alpha+beta) = sinalphacosbeta+sinbetacosalpha (1) sin(alpha-beta) = sinalphacosbeta- sinbetacosalpha (2) cos(alpha+beta) = cosalphacosbeta-sinalphasinbeta (3) cos(alpha-beta) = cosalphacosbeta+sinalphasinbeta ok so we have sin θ cos 3θ + cos θ sin 3θ = 0 sinθ(cos(θ)cos(2θ)-sin(θ)sin(2θ)) + cos(θ)(sin(θ)cos(2θ)+sin(2θ)cos(θ)) = 0

OpenStudy (australopithecus):

sin(3θ) = sin(2θ + θ)

OpenStudy (australopithecus):

sinθ(cos(θ)cos(2θ)-sin(θ)sin(2θ)) + cos(θ)(sin(θ)cos(2θ)+sin(2θ)cos(θ)) = 0 sinθcos(θ)cos(2θ) - sin^(2)(θ)sin(2θ) + cos(θ)sin(θ)cos(2θ) + sin(2θ)cos^(2)(θ) = 0 - sin^(2)(θ)sin(2θ) + sin(2θ)cos^(2)(θ) = 0

OpenStudy (australopithecus):

- sin^(2)(θ)(sin(θ)cos(θ)+sin(θ)cos(θ))+ (sin(θ)cos(θ)+sin(θ)cos(θ))cos^(2)(θ) = 0 -sin^(2)θsin(θ)cos(θ) - sin^(2)(θ)cos(θ) + sin(θ)cos(θ)cos^(2)(θ) + cos^(2)(θ)sin(θ)cos(θ) = 0

OpenStudy (australopithecus):

-sin^(2)θsin(θ)cos(θ) - sin^(2)(θ)cos(θ) + sin(θ)cos(θ)cos^(2)(θ) + cos^(2)(θ)sin(θ)cos(θ) = 0 sin(θ)cos(θ)(-sin^(2)(θ) - sin(θ) + cos^(2)(θ) + cos^(2)(θ)) = 0

OpenStudy (australopithecus):

-sin^(2)(θ) - sin(θ) + cos^(2)(θ) + cos^(2)(θ) = 0 simplify using identites -1+cos^(2)(θ) - sin(θ) +cos^(2)(θ) + cos^(2)(θ) = 0 3cos^(2)(θ) - sin(θ) = 0

OpenStudy (australopithecus):

actually probably best to convert things interms of sin(θ) so I can factor

OpenStudy (australopithecus):

-sin^(2)(θ) - sin(θ) + 1 - sin^(2)(θ) + 1 - sin^(2)(θ) = 0 -3sin^(2)(θ) - sin(θ) + 2 = 0

OpenStudy (anonymous):

You don't need to factor at all. You only need to use the addition or subtraction formula.

OpenStudy (anonymous):

@Australopithecus Actually, you can directly apply the formula \[sin(\alpha+\beta)=sin\alpha cos\beta +sin\beta cos\alpha\] Put α = θ and β =3θ , you can see the the question is actually the same as the right hand side. So, you can get sin (θ+3θ) = sin4θ immediately.

OpenStudy (anonymous):

I know that the answer Sin4Theta is correct, but for some reason Web Assign (my homework site) marks it incorrectly.

OpenStudy (australopithecus):

Email your instructor

OpenStudy (anonymous):

That would be a tech problem, not a math problem anymore!

OpenStudy (anonymous):

(n /π)/4 I've also tried that, but it's marked as incorrect as well.

OpenStudy (anonymous):

I hope it is a tech problem, I've already emailed my instructor. Thanks anyways.

OpenStudy (australopithecus):

the answer is pin/4 where n is an element of Integers you are correct wolfram alpha confirms email instructor